OpenStudy (anonymous):

Evaluate the integral. int_(1)^(16) (x-1)/sqrt(x) dx

Parth (parthkohli):

\[\int_{1}^{16}\dfrac{x-1}{\sqrt{x}}dx\]You can write the integrand as\[=\dfrac{x}{\sqrt x} -\dfrac{1}{\sqrt x}\]

Parth (parthkohli):

\[= x^{1/2} - x^{-1/2}\]Use the power rule.

OpenStudy (kainui):

Another way that's really not better than ParthKohli's way is to choose u=sqrt(x).